The perimeter of a rectangular garden is 80 meters.If its width is 2/3 of its length, find the dimensions of the garden. Represent the data, write the equation , and solve

Answer by algebrahouse.com(1659)

x = length
(2/3)x = width

Perimeter of a rectangle is 2(length) + 2(width)

2x + 2(2/3)x = 80 {perimeter is 2(length) + 2(width)
2x + (4/3)x = 80 {multiplied}
(10/3)x = 80 {combined like terms}
x = 24 {divided each side by 10/3}
(2/3)x = 16 {substituted 24, in for x, into (2/3)x}

length is 24
width is 16

For more help from me, visit:
www.algebrahouse.com